Circular No. 7808

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN OPHIUCHUS Independent discoveries of a possible nova in Ophiuchus have been re-

ported by Katsumi Haseda (Aichi, Japan, 0.10-m f/4 twin patrol camera + T-Max 400 film) and by Yuji Nakamura (Suzuka, Mie, Japan, 200-mm f/4 lens + Tri-X film; communicated by M. Soma, National Astronomical Observatory, Tokyo). Haseda provides the following precise position from his discovery film (Jan. 24.838 UT, when the new object was at mag 9.0), measured by K. Kadota, Ageo, Saitama, Japan: α = 17^{h}37^{m}34^{s}.41, δ = −16^{o}23′17′′.9 (equinox 2000.0). Nakamura provides the following position end figures for the new object (from his discovery film), which was at mag 9.3 on Jan. 24.867: 34^{s}.8, 27′′ (uncertainty ± 20′′ in each coordinate). D. West, Mulvane, KS, reports the following position end figures for the new object, which was at V = 9.19 on a CCD image taken on Jan. 26.527: 34^{s}.60, 23′′.2 (uncertainty ± 0′′.6). Haseda notes that no object appears at the position of the new star to limiting mag 12.5–13.1 on his patrol films taken during 1997 Aug. 23–2001 Oct. 11. The closest star visible on the Digitized Sky Survey (DSS) appears to be a USNO-A2.0 star at position end figures 34^{s}.483, 26′′.09 (blue mag 16.6, red mag 15.0). West adds that the new star was at R_{c} = 8.67 on a CCD image taken on Jan. 26.531, and he states that a comparison with his images indicates that the new object was not visible on the DSS. Visual magnitude estimate by A. Pearce, Nedlands, W. Australia: Jan. 26.857, 9.1.

SUPERNOVA 2002J IN NGC 3464 E. Cappellaro, Osservatorio Astronomico di Capodimonte, Napoli, re-

ports: “A CCD spectrum (range 370–850 nm) of SN 2002J (cf. IAUC 7800), obtained on Jan. 26.25 UT at the European Southern Observatory at La Silla using the New Technology Telescope (+ EMMI), reveals that the object is a type-Ic supernova, very similar to SN 1994I near maximum (cf. Clocchiatti et al. 1996, Ap.J. 462, 462).”

SUPERNOVA 2002an IN NGC 2575 W. M. Wood-Vasey, University of California at Berkeley and Lawrence

Berkeley National Laboratory, reports that observations at the location of SN 2002an (cf. IAUC 7805), taken with the Oschin 1.2-m Schmidt telescope at Palomar on Jan. 17.54 UT, show nothing to a limiting unfiltered magnitude of 19.0, indicating that SN 2002an was caught quite soon after explosion.

2002 January 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green